### v1

# Small Neutrino Masses from a Decoupled Singlet Scalar Field

Abstract
We propose a unified solution with $Z_{4}$ discrete symmetry for small neutrino masses and stability of dark matter. Our model consists of an inert doublet scalar, a dark singlet scalar $S$, a spurion scalar $\varphi$ and right-handed neutrinos beyond the Standard Model (SM), which transform nontrivially under $Z_{4}$. Once the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of $\varphi$, much below the mass scale of the singlet scalar $S$, a small quartic coupling between SM Higgs and inert doublets, which violates both $Z_{4}$ and lepton number, is generated at tree level, so small neutrino masses are obtained at one-loops. We discuss the consistent parameter space for neutrino masses, direct detection and thermal leptogenesis in the model.
Introduction: Neutrino oscillations ParticleDataGroup:2022pth indicate that neutrinos have nonzero masses about $m_{\nu}\sim 0.05\,{\rm eV}$. Moreover, from neutrinoless double beta decay experiments for Majorana neutrinos KATRIN:2021uub as well as Cosmic Microwave Background anisotropies and Baryon Acoustic Oscillations at cosmological scales Planck:2018vyg , we have set upper bounds on the neutrino masses as $m_{{\bar{\nu}}_{e}}<0.8\,{\rm eV}$ and $\sum_{i}m_{\nu_{i}}<0.12\,{\rm eV}$ at 95% CL, respectively. In order to account for small neutrino masses, we need to extend the Standard Model (SM) by the dimension-5 operators for Majorana neutrino masses with a cutoff scale much larger than the electroweak scale or generate neutrino masses of Dirac type from the Yukawa couplings with light sterile neutrinos much below the electroweak scale Cai:2017jrq .
Seesaw mechanism Minkowski:1977sc ; Yanagida:1979as ; Gell-Mann:1979vob ; Mohapatra:1979ia gives rise to small neutrino masses due to the mixing between active neutrinos and heavy Majorana neutrinos. Alternatively, radiative seesaw scenarios provide small neutrino masses by loop corrections with new particles not far from the electroweak scale, due to small loop suppression factors. In the scotogenic model Ma:2006km , neutrino masses are generated by one-loop corrections with right-handed (RH) neutrinos and an extra doublet scalar, which can be a dark matter candidate Barbieri:2006dq ; LopezHonorez:2006gr . However, there is still a need of very small parameters in obtaining small neutrino masses such as a small mixing quartic coupling in the Higgs sector or small Yukawa couplings for the RH neutrinos.
In this article, we propose a novel model with $Z_{4}$ discrete Peccei-Quinn (PQ) symmetry for the common origin of small neutrino masses and stability of dark matter. In our model, there are one inert doublet scalar, three right-handed neutrinos, a real spurion scalar and a singlet complex scalar field $S$, all transforming nontrivially under $Z_{4}$. The lepton number is respected in the renormalizable interactions due to $Z_{4}$. Then, after the $Z_{4}$ symmetry is broken to $Z_{2}$ due to the VEV of the spurion field, the lepton number is broken, and the lightest neutral dark scalar in the model becomes a dark matter candidate.
Taking the decoupling limit for making the singlet scalar field $S$, we show how a small mass splitting between the neutral components of the inert doublet arises, in turn giving rise to small neutrino masses at one-loops. Our results remain intact for any higher order discrete symmetries or a continuous approximate $U(1)$ PQ symmetry containing $Z_{2}$ as a subgroup or in the cases with more complex singlet scalars and inert doublet scalars. We discuss various phenomenological constraints on the $Z_{4}$ violation from neutrino masses, dark matter direction detection and thermal leptogenesis.
Discrete symmetry and dark matter scalars: We introduce a complex singlet scalar $S$ and one extra electroweak doublet scalar, $H_{2}$, other than the SM Higgs doublet $H_{1}$, and three right-handed (RH) neutrinos, $N_{Ri}$, with $i=1,2,3$, and a spurion real scalar field $\varphi$, beyond the SM. We impose a $Z_{4}$ PQ symmetry for the extra fields such that $\phi_{i}\to e^{iQ_{X}\pi/2}\phi_{i}$ with $\phi_{i}=S,\varphi,H_{1},H_{2},N_{R}$ and $Q_{X}(\phi_{i})=(+1,+2,0,+1,+1)$. On the other hand, the SM fermions are neutral under $Z_{4}$ so the model is different from the flavor-dependent $Z_{4}$ symmetry introduced in Ref. Babu:2013pma . As the subgroup of the $Z_{4}$ symmetry, we can identify $Z_{2}$ parities as $Z_{2}(\phi_{i})=(-1,+1,+1,-1,-1)$. The spurion field $\varphi$ transforms by $\varphi\to-\varphi$ under the $Z_{4}$ symmetry but it is invariant under the $Z_{2}$ subgroup. The spurion field $\varphi$ can be a fundamental scalar field or a composite field due to hidden strong dynamics.
Now we consider the relevant interaction Lagrangian for the model as
${\cal L}_{\rm int}$ $=$ $-\sqrt{2}\kappa S^{\dagger}H^{\dagger}_{1}H_{2}-\sqrt{2}\lambda^{\prime}_{S\varphi}\varphi SH^{\dagger}_{1}H_{2}$ (1)
$-y_{N,ij}{\bar{l}}_{i}{\widetilde{H}}_{2}N_{R,j}-\frac{1}{2}\lambda_{N,ij}\varphi\overline{N^{c}_{R,i}}N_{R,j}+{\rm h.c.}$
with ${\widetilde{H}}_{2}=i\sigma_{2}H^{*}_{2}$. It is remarkable that $\frac{1}{2}\lambda_{5}(H^{\dagger}_{1}H_{2})^{2}$ and its complex conjugate, vanish due to the $Z_{4}$ symmetry, and there is no bare mass terms for right-handed neutrinos. We note that there are extra $Z_{4}$ invariant interaction terms in the potential such as $\mu S^{2},\frac{1}{2}\lambda_{H_{i}\varphi}\varphi^{2}|H_{i}|^{2}$($i=1,2$), etc, but it is sufficient to focus on eq. (1) for our discussion below.
After $\varphi$ gets a VEV by $\langle\varphi\rangle=v_{\varphi}$, $Z_{4}$ is broken to $Z_{2}$ and we obtain the $Z_{2}$ invariant effective interactions and the mass terms for right-handed neutrinos,
${\cal L}_{\rm int,eff}$ $=$ $-\sqrt{2}\kappa S^{\dagger}H^{\dagger}_{1}H_{2}-\sqrt{2}\kappa^{\prime}SH^{\dagger}_{1}H_{2}$ (2)
$-y_{N,ij}{\bar{l}}_{i}{\widetilde{H}}_{2}N_{R,j}-\frac{1}{2}M_{N,i}\overline{N^{c}_{R,i}}N_{R,i}+{\rm h.c.}$
where $\kappa^{\prime}\equiv\lambda^{\prime}_{S\varphi}v_{\varphi}$ is the $Z_{4}$ breaking effective interaction and $M_{N,i}\equiv\lambda_{N,ii}v_{\varphi}$ are the masses for RH neutrinos. Here, we assumed the basis where $\lambda_{N,ij}$ is diagonalized. After electroweak symmetry breaking with $\langle H_{1}\rangle=(0,v_{H})^{T}/\sqrt{2}\neq 0$, there appear mass mixings between dark matter scalars, namely, $S$ and the neutral components of $H_{2}$, so nonzero neutrino masses can be generated at one-loops, which will be discussed later.
We remark that $S$ and $H_{2}$ do not get VEVs, leaving a $Z_{2}$ symmetry unbroken. $S$, $H_{2}$ and $N_{Ri}$ are odd under the remaining $Z_{2}$, which ensures the stability of the lightest neutral $Z_{2}$-odd particle among them as a dark matter candidate. In the following discussion, in order to explain small neutrino masses without introducing small parameters, we consider the case where the singlet scalar $S$ is decoupled. In this case, either the lighter neutral component of $H_{2}$ or the lightest RH neutrino can be a dark matter candidate.
Figure 1: Feynman diagrams contributing to the effective $\lambda_{5,{\rm eff}}$ at tree level.
Decoupled scalar and inert doublet scalar masses: From the effective interactions in eq. (2), we can take the limit of integrating out the singlet scalar $S$ whose squared mass becomes $m^{2}_{S}\gg\kappa\kappa^{\prime}$. As a result, we obtain a renormalizable but $Z_{4}$ breaking effective interaction for two Higgs doublets from $\kappa,\kappa^{\prime}$ terms, at tree level as shown in Fig. 1, in the following form,
${\cal L}_{\rm eff}=-\frac{1}{2}\lambda_{5,{\rm eff}}(H^{\dagger}_{1}H_{2})^{2}+{\rm h.c.}$ (3)
with
$\lambda_{5,{\rm eff}}=-\frac{4\kappa\kappa^{\prime}}{m^{2}_{S}}.$ (4)
Here, $\kappa$ is a $Z_{4}$ invariant dimensionful coupling, so it can be of order $m_{S}$. But, $\kappa^{\prime}$ is a $Z_{4}$ breaking term due to the VEV of the spurion field $\varphi$, so it is naturally suppressed as compared to $m_{S}$, that is, $\kappa^{\prime}=\lambda^{\prime}_{S\varphi}v_{\varphi}\ll m_{S}$. Therefore, the effective $\lambda_{5,{\rm eff}}$ term in eq. (4) is also suppressed as a consequence of the decoupled scalar.
For a small $\lambda_{5}$ term in the effective theory below the threshold of the singlet scalar $S$, we obtain the masses for the components of the inert doublet, $H_{2}=(H^{+},\frac{1}{\sqrt{2}}(H_{0}+iA_{0}))^{T}$, as follows,
$m^{2}_{H^{\pm}}$ $=m^{2}_{2}+\frac{1}{2}\lambda_{3}v^{2}_{H},$ (5)
$m^{2}_{H_{0}}$ $=m^{2}_{2}+\frac{1}{2}\left(\lambda_{L}+\lambda_{5,{\rm eff}}\right)v^{2}_{H},$ (6)
$m^{2}_{A_{0}}$ $=m^{2}_{2}+\frac{1}{2}\left(\lambda_{L}-\lambda_{5,{\rm eff}}\right)v^{2}_{H}$ (7)
where $\lambda_{L}\equiv\lambda_{3}+\lambda_{4}$, with $\lambda_{3},\lambda_{4}$ being mixing quartic couplings in the scalar potential, $\lambda_{3}|H_{1}|^{2}|H_{2}|^{2}+\lambda_{4}(H^{\dagger}_{1}H_{2})(H^{\dagger}_{2}H_{1})$, and $m^{2}_{2}$ is the squared mass parameter for the inert doublet $H_{2}$. Henceforth, we choose $\lambda_{5,{\rm eff}}>0$ without loss of generality such that $A_{0}$ is a dark matter candidate.
In the decoupling limit of the singlet scalar $S$, the small effective $\lambda_{5}$ term gives rise to a mass splitting between two real neutral scalars of $H_{2}$ is suppressed as follows, $\Delta\equiv m_{H_{0}}-m_{A_{0}}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{2m_{0}},$ with $m^{2}_{0}\equiv(m^{2}_{H_{0}}+m^{2}_{A_{0}})/2$. The lighter neutral component of the inert doublet $H_{2}$ is a dark matter candidate avoiding the strong bound on DM-nucleon scattering with $Z$ boson exchange, as far as $|\Delta|\gtrsim 100\,{\rm keV}$, which sets a lower bound on $\lambda_{5,{\rm eff}}$ as $\lambda_{5,{\rm eff}}\gtrsim 3.3\times 10^{-7}(m_{0}/100\,{\rm GeV})(|\Delta|/100\,{\rm keV})$. Then, for $\kappa\sim m_{S}$, we need $\kappa^{\prime}/m_{S}\gtrsim 8\times 10^{-8}$ for $m_{0}\sim 100\,{\rm GeV}$ and $|\Delta|\gtrsim 100\,{\rm keV}$.
Figure 2: Feynman diagrams contributing to neutrino masses at one-loop.
Decoupled scalar and small neutrino masses: In the decoupling limit of the singlet scalar $S$, from the combination of the Yukawa interactions for right-handed neutrinos and a small effective $\lambda_{5,{\rm eff}}$ term in eq. (3), we obtain the one-loop corrections for neutrino masses from the Feynman diagram shown in Fig. 2 as
$({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}}\sum_{k}\frac{y_{N,ik}y_{N,jk}M_{N,k}}{m^{2}_{0}-M^{2}_{N,k}}\bigg{[}1-\frac{M^{2}_{N,k}}{m^{2}_{0}-M^{2}_{N,k}}\ln\frac{m^{2}_{0}}{M^{2}_{N,k}}\bigg{]}.$ (8)
The above results match with those due to a nonzero small $\lambda_{5}$ in the case without a $Z_{4}$ symmetry in Ref. Ma:2006km . Then, when the RH neutrinos are heavier than the inert doublet Higgs, namely, $M_{N,k}\gg m_{0}$, the above expression with eq. (4) gets simplified to
$({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}}\sum_{k}\frac{y_{N,ik}y_{N,jk}}{M_{N,k}}\bigg{[}\ln\frac{M^{2}_{N,k}}{m^{2}_{0}}-1\bigg{]}.$ (9)
Therefore, for neutrino masses, $m_{\nu}\lesssim 0.1\,{\rm eV}$, choosing $y_{N,ij}={\cal O}(1)$ and $M_{N,k}\sim 10^{9}\,{\rm GeV}$, we can set the upper bound on $\lambda_{5,{\rm eff}}$ as $\lambda_{5,{\rm eff}}\lesssim 8\times 10^{-6}$, which is consistent with the lower bound on $\Delta$ from the direct detection. In other words, we need $\kappa^{\prime}/m_{S}\lesssim 10^{-6}$ for $\kappa\sim m_{S}$ and $M_{N,k}\lesssim\kappa^{\prime}\ll m_{S}$ for $\lambda_{N,kk}\lesssim\lambda^{\prime}_{S\varphi}$. Then, we can keep the RH right-handed neutrinos in the effective theory while the singlet scalar $S$ is decoupled. So, for a small $\lambda_{5,{\rm eff}}$, it is sufficient to take $m_{S}\gtrsim M_{N,k}\sim 10^{9}\,{\rm GeV}$, unless we tune the Yukawa couplings $y_{N,ij}$ to small values.
Figure 3: Correlations for inert doublet scalar mass $m_{0}$ vs RH neutrino mass $M_{N}$ for varying $\lambda_{5,{\rm eff}}$ (left plot); neutrino mass $m_{\nu}$ vs singlet scalar mass $m_{S}$ for varying $M_{N}$ (right plot).
We also comment on the case with $m_{0}\gg M_{N,k}$. In this case, the lightest RH neutrino is a dark matter candidate Ibarra:2016dlb whereas the scalars of the inert doublet decay into a pair of the SM lepton and the RH neutrino through the Yukawa interactions. In this limit, eq. (8) becomes approximated to $({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}m^{2}_{0}}\sum_{k}y_{N,ik}y_{N,jk}M_{N,k}.$ For instance, for $m_{\nu}\lesssim 0.1\,{\rm eV}$, $y_{N,ij}={\cal O}(1)$, $m_{0}\sim 100\,{\rm GeV}$ and $M_{N,k}\sim 10\,{\rm MeV}$, we can set the upper bound, $\lambda_{5,{\rm eff}}\lesssim 8\times 10^{-10}$ or $|\Delta|\lesssim 0.2\,{\rm keV}$, but there is no more lower bound on $\lambda_{5,{\rm eff}}$ because the scalars in the inert doublet decay. A small effective $\lambda_{5,{\rm eff}}$ is achievable if $m_{S}\gtrsim 10^{9}\,{\rm GeV}(\lambda^{\prime}_{S\varphi}/\lambda_{N,kk})$ for $v_{\varphi}=10\,{\rm MeV}/\lambda_{N,kk}$ for $\kappa\sim m_{S}$.
In the left plot of Fig. 3, we show the parameter space for inert doublet scalar mass $m_{0}$ vs RH neutrino mass $M_{N}\equiv M_{N,1}$ satisfying $m_{\nu}=0.1\,{\rm eV}$. The color code indicates a scan over $\lambda_{5,{\rm eff}}$ between $10^{-8}$ and $10^{-4}$, showing that $\lambda_{5,{\rm eff}}\gtrsim 10^{-6}$ and $M_{N}\gtrsim 2\times 10^{7}\,{\rm GeV}$ are necessary for $m_{0}\gtrsim 100\,{\rm GeV}$. In the right plot of Fig. 3, we decode the correlation between neutrino mass $m_{\nu}$ and singlet scalar mass $m_{S}$ for varying $M_{N}$ between $10^{4}\,{\rm GeV}$ and $10^{9}\,{\rm GeV}$. In both plots in Fig. 3, we imposed $|\Delta|>100\,{\rm keV}$ and took into account the perturbativity, $|\lambda_{5,{\rm eff}}|<1$. Then, we need the singlet scalar mass $m_{S}$ to be of order $10^{14}-10^{16}\,{\rm GeV}$ for $m_{\nu}=0.01-0.1\,{\rm eV}$, which might hint at a connection of neutrino mass generation to the unification. If we choose smaller Yukawa couplings $y_{N}$, the diagonal boundary in the left plot of Fig. 3 is shifted to the left, allowing for smaller $M_{N}$ for a fixed $m_{0}$, and the strip in the right plot of Fig. 3 is shifted to the bottom, making $m_{S}$ smaller for a fixed $m_{\nu}$.
Dark matter direct detection: When the lighter neutral scalar from the inert doublet is dark matter, we can avoid the strong bound from direct detection if the mass splitting between $A_{0}$ and $H_{0}$ is large enough, i.e. $|\Delta|\gtrsim 100\,{\rm keV}$. Then, the cross section for spin-independent scattering between dark matter and nucleus is dominated by the tree-level Higgs exchange Barbieri:2006dq ; LopezHonorez:2006gr , as follows,
$\sigma^{\rm SI}_{A_{0}-N}=\frac{(\lambda_{L}-\lambda_{5,\,{\rm eff}})^{2}\mu^{2}_{N}}{4\pi m^{2}_{A_{0}}m^{4}_{h}A^{2}}(Zf_{p}+(A-Z)f_{n})^{2}$ (10)
where $Z,A-Z$ are the numbers of protons and neutrons in the detector nucleus, $\mu_{N}$ is the reduced mass for the DM-nucleus system, and $f_{p,n}$ are the nucleon form factors Hoferichter:2015dsa ; Junnarkar:2013ac ; Kim:2023pwf .
The DM relic density is determined by the DM annihilations into a pair of SM fermions, $WW,ZZ$ and $h_{\rm SM}h_{\rm SM}$ with $h_{\rm SM}$ being the SM Higgs boson Barbieri:2006dq ; LopezHonorez:2006gr ; Gong:2012ri . In particular, there is a room to adjust $\lambda_{L}$ to explain the DM relic density for relatively heavy DM due to $A_{0}A_{0}\to h_{\rm SM}h_{\rm SM}$ while the direct detection bound is satisfied Belyaev:2016lok . DM masses above about $100\,{\rm GeV}$ are consistent with the bounds from direct detection, as far as $|\lambda_{L}-\lambda_{5,\,{\rm eff}}|\lesssim 10^{-2}$, apart from the resonance. But, if $|\lambda_{L}-\lambda_{5,\,{\rm eff}}|\lesssim 10^{-3}$, as required near resonances, the one-loop corrections to the DM-nucleus cross section are relevant Abe:2015rja .
Thermal leptogenesis: Baryogenesis can be realized in our model via thermal leptogenesis Fukugita:1986hr , due to the decays of RH neutrinos with hierarchical masses. As the asymmetries generated from the decays of the heavier RH neutrinos, $N_{2,3}$, are negligible due to strong washout effects by scattering, we focus on the $N_{1}$ contribution to the lepton asymmetry. The baryon-to-photon ratio converted from the final $B-L$ asymmetry is given by
$\eta_{B}=-C\varepsilon_{1}\kappa_{1}$ (11)
where $C\simeq 0.01$, $\varepsilon_{1}$ is the $CP$ asymmetry parameter, and $\kappa_{1}$ is the efficiency factor approximated Buchmuller:2004nz to
$\kappa_{1}(K_{1})\simeq\frac{2}{z_{B}(K_{1})K_{1}}\Big{(}1-e^{-\frac{1}{2}z_{B}(K_{1})K_{1}}\Big{)},$ (12)
with $K_{1}=\Gamma_{1}/H(z_{1}=1)$ being the decay parameter and $z_{B}$ being the value of $z_{1}=M_{N,1}/T$ at the completion of the baryon asymmetry generation, given Buchmuller:2004nz by
$z_{B}(K)\simeq 1+\frac{1}{2}\ln\bigg{(}1+\frac{\pi K^{2}}{1024}\bigg{[}\ln\Big{(}\frac{3125\pi K^{2}}{1024}\Big{)}\bigg{]}^{5}\bigg{)}.$ (13)
Here, $\Gamma_{1}$ is the decay rate for $N_{1}$, given by $\Gamma_{1}\simeq\frac{M_{N,1}}{8\pi}(y^{\dagger}_{N}y_{N})_{11}$ for $m_{0}\ll M_{N,1}$, and $H(z_{1})$ is the Hubble parameter. In the Casas-Ibarra parametrization for the Yukawa matrix $y_{N}$ Casas:2001sr , we can express $(y^{\dagger}_{N}y_{N})_{ij}=\sqrt{\Lambda_{i}\Lambda_{j}}{\widetilde{m}}_{ij}$ in terms of the effective neutrino mass matrix ${\widetilde{m}}\equiv RD_{{\cal M}_{\nu}}R^{\dagger}$ with $R$ being arbitrary complex rotation matrix, $D_{{\cal M}_{\nu}}={\rm diag}(m_{\nu_{1}},m_{\nu_{2}},m_{\nu_{3}})$ and $\Lambda_{i}=\frac{4\pi^{2}\xi_{i}M_{N,i}}{\lambda_{5,{\rm eff}}v^{2}_{H}}$ Hugle:2018qbw . Here, $\xi_{i}=\big{(}\frac{1}{8}\frac{M^{2}_{N,i}}{m^{2}_{H_{0}}-m^{2}_{A_{0}}}[L(m^{2}_{H_{0}})-L(m^{2}_{A_{0}})]\big{)}^{-1}$ with $L(m^{2})=\frac{m^{2}}{m^{2}-M^{2}_{N,i}}\ln(m^{2}/M^{2}_{N,i})$ are order one parameters in most of the parameter space. Thus, we get the decay parameter as $K_{1}\simeq\frac{2\pi^{2}\xi_{1}}{\lambda_{5\,{\rm eff}}}\sqrt{\frac{45}{64\pi^{5}g_{*}}}\frac{M_{P}}{v^{2}_{H}}{\widetilde{m}}_{11}$ for $m_{0}\ll M_{N,1}$.
For the case with three RH neutrinos, the $CP$ asymmetry parameter is bounded Davidson:2002qv ; Hugle:2018qbw by
$\varepsilon_{1}\lesssim\frac{3\pi}{4\lambda_{5\,{\rm eff}}v^{2}_{H}}\,\xi_{3}(m_{h}-m_{l})M_{N,1}$ (14)
with $m_{h},m_{l}$ are the masses for the heaviest and lightest active neutrinos, respectively. After optimizing the efficiency factor $\kappa_{1}$ and the $CP$ asymmetry parameter $\varepsilon_{1}$, we obtain the relation, $\lambda_{5,{\rm eff}}\simeq 4\pi\xi_{1}(m_{l}/10^{-3}\,{\rm eV})$ Hugle:2018qbw , for which the baryon-to-photon ratio without washout effects from $\Delta L$ processes is given by
$\eta_{B}\simeq 3.8\times 10^{-10}\Big{(}\frac{M_{N,1}}{10^{4}\,{\rm GeV}}\Big{)}\bigg{(}\frac{10^{-4}}{\lambda_{5,{\rm eff}}}\bigg{)}\bigg{(}\frac{m_{h}}{0.1\,{\rm eV}}\bigg{)}.$ (15)
Figure 4: Parameter space for RH neutrino mass $M_{N}$ vs $\lambda_{5,{\rm eff}}$, showing $\eta_{B}>6.1\times 10^{-10}$ in purple region.
[…]
We remark that the asymmetry in the inert doublet or the DM asymmetry can reduce the lepton asymmetry, because $lH_{2}\to{\bar{l}}{\bar{H}}_{2}$ is favored over ${\bar{l}}{\bar{H}_{2}}\to lH_{2}$ Racker:2014yfa . However, for a sizable $\lambda_{5,{\rm eff}}$, $H_{2}{\bar{H}}_{1}\leftrightarrow{\bar{H}}_{2}H_{1}$ process is fast, so it can make the DM asymmetry Boltzmann sufficiently suppressed. It turns out that $\Delta L=2$ processes are inefficient if $|\lambda_{5,{\rm eff}}|\gtrsim 10^{-5}$ for $M_{N,1}\lesssim 10^{9}\,{\rm GeV}$ Clarke:2015hta .
Conclusions: We presented a unified solution with $Z_{4}$ discrete Peccei-Quinn symmetry for small neutrino masses and stability of dark matter in the radiative seesaw scenarios. Once the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of an extra singlet scalar field $\varphi$ much below the mass scale of the singlet scalar $S$, a small mixing between the dark sector neutral scalars is generated. Then, we showed that small neutrino masses are obtained from one-loop diagrams with RH neutrinos and inert doublet scalars with small mass splittings and identified the parameter space where observed neutrino masses are explained and the lighter scalar of the inert doublet becomes a dark matter candidate. From the measurement of the absolute neutrino masses at the level of $0.01\,{\rm eV}$, direct detection searches for the inert doublet dark matter and the successful leptogenesis, we showed that a new mass scale of the singlet scalar $S$ is expected to be up to $\sim 10^{16}\,{\rm GeV}$.
Acknowledgments
The work is supported in part by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567 and RS-2024-00341419 (JKK)).

### v2

@@
 We propose a unified solution with $Z_{4}$ discrete symmetry for small neutrino masses and stability of dark matter.
-Our model consists of an inert doublet scalar, a dark singlet scalar $S$, a spurion scalar $\varphi$ and right-handed neutrinos beyond the Standard Model (SM), which transform nontrivially under $Z_{4}$.
-Once the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of $\varphi$, much below the mass scale of the singlet scalar $S$, a small quartic coupling between SM Higgs and inert doublets, which violates both $Z_{4}$ and lepton number, is generated at tree level, so small neutrino masses are obtained at one-loops.
-We discuss the consistent parameter space for neutrino masses, direct detection and thermal leptogenesis in the model.
-Introduction: Neutrino oscillations ParticleDataGroup:2022pth indicate that neutrinos have nonzero masses about $m_{\nu}\sim 0.05\,{\rm eV}$.
-Moreover, from neutrinoless double beta decay experiments for Majorana neutrinos KATRIN:2021uub as well as Cosmic Microwave Background anisotropies and Baryon Acoustic Oscillations at cosmological scales Planck:2018vyg , we have set upper bounds on the neutrino masses as $m_{{\bar{\nu}}_{e}}<0.8\,{\rm eV}$ and $\sum_{i}m_{\nu_{i}}<0.12\,{\rm eV}$ at 95% CL, respectively.
-In order to account for small neutrino masses, we need to extend the Standard Model (SM) by the dimension-5 operators for Majorana neutrino masses with a cutoff scale much larger than the electroweak scale or generate neutrino masses of Dirac type from the Yukawa couplings with light sterile neutrinos much below the electroweak scale Cai:2017jrq .
-Seesaw mechanism Minkowski:1977sc ; Yanagida:1979as ; Gell-Mann:1979vob ; Mohapatra:1979ia gives rise to small neutrino masses due to the mixing between active neutrinos and heavy Majorana neutrinos.
-Alternatively, radiative seesaw scenarios provide small neutrino masses by loop corrections with new particles not far from the electroweak scale, due to small loop suppression factors.
-In the scotogenic model Ma:2006km , neutrino masses are generated by one-loop corrections with right-handed (RH) neutrinos and an extra doublet scalar, which can be a dark matter candidate Barbieri:2006dq ; LopezHonorez:2006gr .
-However, there is still a need of very small parameters in obtaining small neutrino masses such as a small mixing quartic coupling in the Higgs sector or small Yukawa couplings for the RH neutrinos.
-In this article, we propose a novel model with $Z_{4}$ discrete Peccei-Quinn (PQ) symmetry for the common origin of small neutrino masses and stability of dark matter.
-In our model, there are one inert doublet scalar, three right-handed neutrinos, a real spurion scalar and a singlet complex scalar field $S$, all transforming nontrivially under $Z_{4}$.
+The Standard Model is extended with an inert doublet scalar, a dark singlet scalar, a spurion scalar and right-handed neutrinos, which all transform nontrivially under $Z_{4}$.
+After the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion, much below the mass scale of the dark singlet scalar, a small lepton number violating coupling for the inert doublet is generated at tree level, so small neutrino masses are obtained at one-loops.
+We discuss the important roles of the $Z_{4}$ symmetry for neutrino masses, dark matter physics and thermal leptogenesis.
+I Introduction
+Neutrino oscillations ParticleDataGroup:2022pth indicate neutrinos masses of $m_{\nu}\sim 0.05\,{\rm eV}$.
+Moreover, the bounds from neutrinoless double beta decay experiments KATRIN:2021uub and Cosmic Microwave Background anisotropies and Baryon Acoustic Oscillations Planck:2018vyg are $m_{{\bar{\nu}}_{e}}<0.8\,{\rm eV}$ and $\sum_{i}m_{\nu_{i}}<0.12\,{\rm eV}$ at 95% CL, respectively.
+In order to account for small neutrino masses, it is possible to extend the Standard Model (SM) by the dimension-5 operators for neutrino masses with a cutoff scale beyond the electroweak scale or light sterile neutrinos coupled to neutrinos much below the electroweak scale Cai:2017jrq .
+Small neutrino masses can be obtained by the mixing between active neutrinos and heavy Majorana neutrinos Minkowski:1977sc ; Yanagida:1979as ; Gell-Mann:1979vob ; Mohapatra:1979ia .
+Alternatively, small neutrino masses can be generated by loop corrections, for instance, due to right-handed (RH) neutrinos and an extra doublet scalar Ma:2006km , the latter of which serves as a dark matter candidate Barbieri:2006dq ; LopezHonorez:2006gr .
+However, there is still a need of very small parameters in obtaining small neutrino masses in similar models Cai:2017jrq .
+In this article, we propose a novel model with a $Z_{4}$ discrete symmetry for the common origin of small neutrino masses and stability of dark matter.
+In our model, there are one inert doublet scalar, three right-handed neutrinos, a real spurion scalar and a singlet complex scalar field, all transforming nontrivially under $Z_{4}$.
 The lepton number is respected in the renormalizable interactions due to $Z_{4}$.
 Then, after the $Z_{4}$ symmetry is broken to $Z_{2}$ due to the VEV of the spurion field, the lepton number is broken, and the lightest neutral dark scalar in the model becomes a dark matter candidate.
-Taking the decoupling limit for making the singlet scalar field $S$, we show how a small mass splitting between the neutral components of the inert doublet arises, in turn giving rise to small neutrino masses at one-loops.
-Our results remain intact for any higher order discrete symmetries or a continuous approximate $U(1)$ PQ symmetry containing $Z_{2}$ as a subgroup or in the cases with more complex singlet scalars and inert doublet scalars.
-We discuss various phenomenological constraints on the $Z_{4}$ violation from neutrino masses, dark matter direction detection and thermal leptogenesis.
-Discrete symmetry and dark matter scalars: We introduce a complex singlet scalar $S$ and one extra electroweak doublet scalar, $H_{2}$, other than the SM Higgs doublet $H_{1}$, and three right-handed (RH) neutrinos, $N_{Ri}$, with $i=1,2,3$, and a spurion real scalar field $\varphi$, beyond the SM.
-We impose a $Z_{4}$ PQ symmetry for the extra fields such that $\phi_{i}\to e^{iQ_{X}\pi/2}\phi_{i}$ with $\phi_{i}=S,\varphi,H_{1},H_{2},N_{R}$ and $Q_{X}(\phi_{i})=(+1,+2,0,+1,+1)$.
-On the other hand, the SM fermions are neutral under $Z_{4}$ so the model is different from the flavor-dependent $Z_{4}$ symmetry introduced in Ref.
+The spurion scalar can be regarded as a pseudo-Goldstone boson coming from the breaking of a global $U(1)$ symmetry into $Z_{4}$ at a higher scale, and $Z_{4}$ can be further down to $Z_{2}$ at a low scale.
+Taking the decoupling limit for making the singlet scalar field $S$ heavy, we show how a small mass splitting between the neutral components of the inert doublet arises, in turn giving rise to small neutrino masses at one-loops.
+We discuss the roles of the new singlet scalars and constraints for neutrino masses, dark matter and thermal leptogenesis.
+II Discrete symmetry and dark matter scalars
+We introduce a complex singlet scalar $S$ and one extra electroweak doublet scalar, $H_{2}$, other than the SM Higgs doublet $H_{1}$, and three right-handed (RH) neutrinos, $N_{R,i}$, with $i=1,2,3$, and a spurion real scalar field $\varphi$, beyond the SM.
+We impose a $Z_{4}$ symmetry for the extra fields such that $\phi_{i}\to e^{iQ_{X}\pi/2}\phi_{i}$ with $\phi_{i}=S,\varphi,H_{1},H_{2},N_{R}$ and $Q_{X}(\phi_{i})=(+1,+2,0,+1,+1)$.
+But, the SM fermions are neutral under $Z_{4}$ so the model is different from the flavor-dependent $Z_{4}$ symmetry introduced in Ref.
 Babu:2013pma .
-As the subgroup of the $Z_{4}$ symmetry, we can identify $Z_{2}$ parities as $Z_{2}(\phi_{i})=(-1,+1,+1,-1,-1)$.
+The $Z_{4}$ symmetry can be embedded into a global $U(1)$ Peccei-Quinn symmetry which is anomalous in the presence of the couplings to the SM quarks or extra quarks, but we don’t pursue this possibility in detail in the current work.
+As the subgroup of the $Z_{4}$ symmetry, we can identify $Z_{2}$ parities as $Z_{2}(\phi_{i})=(-,+,+,-,-)$.
 The spurion field $\varphi$ transforms by $\varphi\to-\varphi$ under the $Z_{4}$ symmetry but it is invariant under the $Z_{2}$ subgroup.
-The spurion field $\varphi$ can be a fundamental scalar field or a composite field due to hidden strong dynamics.
 Now we consider the relevant interaction Lagrangian for the model as
@@
 with ${\widetilde{H}}_{2}=i\sigma_{2}H^{*}_{2}$.
-It is remarkable that $\frac{1}{2}\lambda_{5}(H^{\dagger}_{1}H_{2})^{2}$ and its complex conjugate, vanish due to the $Z_{4}$ symmetry, and there is no bare mass terms for right-handed neutrinos.
-We note that there are extra $Z_{4}$ invariant interaction terms in the potential such as $\mu S^{2},\frac{1}{2}\lambda_{H_{i}\varphi}\varphi^{2}|H_{i}|^{2}$($i=1,2$), etc, but it is sufficient to focus on eq.
+It is remarkable that $\frac{1}{2}\lambda_{5}(H^{\dagger}_{1}H_{2})^{2}$ and its hermitian conjugate, vanish due to the $Z_{4}$ symmetry, and there is no bare mass terms for right-handed neutrinos due to the lepton number conservation.
+There are extra $Z_{4}$ invariant interaction terms in the potential such as $\mu\varphi S^{2},\frac{1}{2}\lambda_{H_{i}\varphi}\varphi^{2}|H_{i}|^{2}$($i=1,2$), etc, as shown in the appendix, but it is sufficient to focus on eq.
 (1) for our discussion below.
@@
 After electroweak symmetry breaking with $\langle H_{1}\rangle=(0,v_{H})^{T}/\sqrt{2}\neq 0$, there appear mass mixings between dark matter scalars, namely, $S$ and the neutral components of $H_{2}$, so nonzero neutrino masses can be generated at one-loops, which will be discussed later.
+When the SM Higgs and the spurion field $\varphi$ are expanded around the VEVs as $H_{1}=(0,v_{H}+h)^{T}/\sqrt{2}$ and $\varphi=v_{\varphi}+\rho$, the dark Higgs $\rho$ can mix with the SM Higgs $h$, so the mass eigenstates become $h_{1}=c_{\alpha}\,h+s_{\alpha}\,\rho$ and $h_{2}=-s_{\alpha}\,h+c_{\alpha}\,\rho$, with $c_{\alpha}\equiv\cos\alpha,s_{\alpha}\equiv\sin\alpha$.
 We remark that $S$ and $H_{2}$ do not get VEVs, leaving a $Z_{2}$ symmetry unbroken.
-$S$, $H_{2}$ and $N_{Ri}$ are odd under the remaining $Z_{2}$, which ensures the stability of the lightest neutral $Z_{2}$-odd particle among them as a dark matter candidate.
-In the following discussion, in order to explain small neutrino masses without introducing small parameters, we consider the case where the singlet scalar $S$ is decoupled.
-In this case, either the lighter neutral component of $H_{2}$ or the lightest RH neutrino can be a dark matter candidate.
-Figure 1: Feynman diagrams contributing to the effective $\lambda_{5,{\rm eff}}$ at tree level.
-Decoupled scalar and inert doublet scalar masses: From the effective interactions in eq.
+$S$, $H_{2}$ and $N_{R,i}$ are odd under the remaining $Z_{2}$, which ensures the stability of the lightest neutral $Z_{2}$-odd particle among them as a dark matter candidate.
+For small neutrino masses without introducing small parameters, we consider the case where the singlet scalar $S$ is decoupled, so either the lighter neutral component of $H_{2}$ or the lightest RH neutrino can be a dark matter candidate.
+III Decoupled scalar and inert doublet scalar masses
+From the effective interactions in eq.
 (2), we can take the limit of integrating out the singlet scalar $S$ whose squared mass becomes $m^{2}_{S}\gg\kappa\kappa^{\prime}$.
@@
 Here, $\kappa$ is a $Z_{4}$ invariant dimensionful coupling, so it can be of order $m_{S}$.
-But, $\kappa^{\prime}$ is a $Z_{4}$ breaking term due to the VEV of the spurion field $\varphi$, so it is naturally suppressed as compared to $m_{S}$, that is, $\kappa^{\prime}=\lambda^{\prime}_{S\varphi}v_{\varphi}\ll m_{S}$.
+But, $\kappa^{\prime}$ is a $Z_{4}$ breaking coupling due to the VEV of the spurion field $\varphi$, so it is naturally suppressed as compared to $m_{S}$, that is, $\kappa^{\prime}=\lambda^{\prime}_{S\varphi}v_{\varphi}\ll m_{S}$.
 Therefore, the effective $\lambda_{5,{\rm eff}}$ term in eq.
 (4) is also suppressed as a consequence of the decoupled scalar.
+Figure 1: Feynman diagrams contributing to $\lambda_{5,{\rm eff}}$.
 For a small $\lambda_{5}$ term in the effective theory below the threshold of the singlet scalar $S$, we obtain the masses for the components of the inert doublet, $H_{2}=(H^{+},\frac{1}{\sqrt{2}}(H_{0}+iA_{0}))^{T}$, as follows,
@@
 Henceforth, we choose $\lambda_{5,{\rm eff}}>0$ without loss of generality such that $A_{0}$ is a dark matter candidate.
-In the decoupling limit of the singlet scalar $S$, the small effective $\lambda_{5}$ term gives rise to a mass splitting between two real neutral scalars of $H_{2}$ is suppressed as follows, $\Delta\equiv m_{H_{0}}-m_{A_{0}}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{2m_{0}},$ with $m^{2}_{0}\equiv(m^{2}_{H_{0}}+m^{2}_{A_{0}})/2$.
-The lighter neutral component of the inert doublet $H_{2}$ is a dark matter candidate avoiding the strong bound on DM-nucleon scattering with $Z$ boson exchange, as far as $|\Delta|\gtrsim 100\,{\rm keV}$, which sets a lower bound on $\lambda_{5,{\rm eff}}$ as $\lambda_{5,{\rm eff}}\gtrsim 3.3\times 10^{-7}(m_{0}/100\,{\rm GeV})(|\Delta|/100\,{\rm keV})$.
+The most general results for the mass eigenvalues and the mixing angles with the non-decoupled singlet scalar $S$ are shown in eqs.
+(39), (50) and (51), in the appendix.
+In the decoupling limit of the singlet scalar $S$, the small effective $\lambda_{5}$ term gives rise to a mass splitting between two real neutral scalars of $H_{2}$ is suppressed as $\Delta\equiv m_{H_{0}}-m_{A_{0}}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{2m_{0}},$ with $m^{2}_{0}\equiv(m^{2}_{H_{0}}+m^{2}_{A_{0}})/2$.
+So, we can avoid the strong bound on DM-nucleon scattering with $Z$ boson exchange, as far as $|\Delta|\gtrsim 100\,{\rm keV}$, which sets a lower bound on $\lambda_{5,{\rm eff}}$ as $\lambda_{5,{\rm eff}}\gtrsim 3.3\times 10^{-7}(m_{0}/100\,{\rm GeV})(|\Delta|/100\,{\rm keV})$.
 Then, for $\kappa\sim m_{S}$, we need $\kappa^{\prime}/m_{S}\gtrsim 8\times 10^{-8}$ for $m_{0}\sim 100\,{\rm GeV}$ and $|\Delta|\gtrsim 100\,{\rm keV}$.
 Figure 2: Feynman diagrams contributing to neutrino masses at one-loop.
-Decoupled scalar and small neutrino masses: In the decoupling limit of the singlet scalar $S$, from the combination of the Yukawa interactions for right-handed neutrinos and a small effective $\lambda_{5,{\rm eff}}$ term in eq.
+IV Decoupled scalar and small neutrino masses
+In the decoupling limit of the singlet scalar $S$, from the combination of the Yukawa interactions for right-handed neutrinos and a small effective $\lambda_{5,{\rm eff}}$ term in eq.
 (3), we obtain the one-loop corrections for neutrino masses from the Feynman diagram shown in Fig.
 2 as
-$({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}}\sum_{k}\frac{y_{N,ik}y_{N,jk}M_{N,k}}{m^{2}_{0}-M^{2}_{N,k}}\bigg{[}1-\frac{M^{2}_{N,k}}{m^{2}_{0}-M^{2}_{N,k}}\ln\frac{m^{2}_{0}}{M^{2}_{N,k}}\bigg{]}.$ (8)
+$({\cal M}_{\nu})_{ij}$ $\simeq$ $\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}}\sum_{k}\frac{y_{N,ik}y_{N,jk}M_{N,k}}{m^{2}_{0}-M^{2}_{N,k}}$ (8)
+×$\times\bigg{[}1-\frac{M^{2}_{N,k}}{m^{2}_{0}-M^{2}_{N,k}}\ln\frac{m^{2}_{0}}{M^{2}_{N,k}}\bigg{]}.$
 The above results match with those due to a nonzero small $\lambda_{5}$ in the case without a $Z_{4}$ symmetry in Ref.
-Ma:2006km .
+Ma:2006km , and the same results can be obtained from the decoupling limit of the general results in eq.
+(53) in the appendix.
 Then, when the RH neutrinos are heavier than the inert doublet Higgs, namely, $M_{N,k}\gg m_{0}$, the above expression with eq.
@@
 So, for a small $\lambda_{5,{\rm eff}}$, it is sufficient to take $m_{S}\gtrsim M_{N,k}\sim 10^{9}\,{\rm GeV}$, unless we tune the Yukawa couplings $y_{N,ij}$ to small values.
-Figure 3: Correlations for inert doublet scalar mass $m_{0}$ vs RH neutrino mass $M_{N}$ for varying $\lambda_{5,{\rm eff}}$ (left plot); neutrino mass $m_{\nu}$ vs singlet scalar mass $m_{S}$ for varying $M_{N}$ (right plot).
-We also comment on the case with $m_{0}\gg M_{N,k}$.
-In this case, the lightest RH neutrino is a dark matter candidate Ibarra:2016dlb whereas the scalars of the inert doublet decay into a pair of the SM lepton and the RH neutrino through the Yukawa interactions.
+In the case with $m_{0}\gg M_{N,k}$, the lightest RH neutrino is a dark matter candidate Ibarra:2016dlb whereas the scalars of the inert doublet decay into a pair of the SM lepton and the RH neutrino through the Yukawa interactions.
 In this limit, eq.
-(8) becomes approximated to $({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}m^{2}_{0}}\sum_{k}y_{N,ik}y_{N,jk}M_{N,k}.$ For instance, for $m_{\nu}\lesssim 0.1\,{\rm eV}$, $y_{N,ij}={\cal O}(1)$, $m_{0}\sim 100\,{\rm GeV}$ and $M_{N,k}\sim 10\,{\rm MeV}$, we can set the upper bound, $\lambda_{5,{\rm eff}}\lesssim 8\times 10^{-10}$ or $|\Delta|\lesssim 0.2\,{\rm keV}$, but there is no more lower bound on $\lambda_{5,{\rm eff}}$ because the scalars in the inert doublet decay.
+(8) becomes approximated to $({\cal M}_{\nu})_{ij}\simeq\frac{\lambda_{5,{\rm eff}}v^{2}_{H}}{16\pi^{2}m^{2}_{0}}\sum_{k}y_{N,ik}y_{N,jk}M_{N,k}.$ For instance, for $m_{\nu}\lesssim 0.1\,{\rm eV}$, $y_{N,ij}={\cal O}(1)$, $m_{0}\sim 100\,{\rm GeV}$ and $M_{N,k}\sim 10\,{\rm MeV}$, we can set the upper bound, $\lambda_{5,{\rm eff}}\lesssim 8\times 10^{-10}$ or $|\Delta|\lesssim 0.2\,{\rm keV}$.
+But, there is no more lower bound on $\lambda_{5,{\rm eff}}$ because the scalars in the inert doublet decay.
 A small effective $\lambda_{5,{\rm eff}}$ is achievable if $m_{S}\gtrsim 10^{9}\,{\rm GeV}(\lambda^{\prime}_{S\varphi}/\lambda_{N,kk})$ for $v_{\varphi}=10\,{\rm MeV}/\lambda_{N,kk}$ for $\kappa\sim m_{S}$.
-In the left plot of Fig.
+In Fig.
 3, we show the parameter space for inert doublet scalar mass $m_{0}$ vs RH neutrino mass $M_{N}\equiv M_{N,1}$ satisfying $m_{\nu}=0.1\,{\rm eV}$.
 The color code indicates a scan over $\lambda_{5,{\rm eff}}$ between $10^{-8}$ and $10^{-4}$, showing that $\lambda_{5,{\rm eff}}\gtrsim 10^{-6}$ and $M_{N}\gtrsim 2\times 10^{7}\,{\rm GeV}$ are necessary for $m_{0}\gtrsim 100\,{\rm GeV}$.
-In the right plot of Fig.
-3, we decode the correlation between neutrino mass $m_{\nu}$ and singlet scalar mass $m_{S}$ for varying $M_{N}$ between $10^{4}\,{\rm GeV}$ and $10^{9}\,{\rm GeV}$.
-In both plots in Fig.
-3, we imposed $|\Delta|>100\,{\rm keV}$ and took into account the perturbativity, $|\lambda_{5,{\rm eff}}|<1$.
+In Fig.
+4, we decode the correlation between neutrino mass $m_{\nu}$ and singlet scalar mass $m_{S}$ for varying $M_{N}$ between $10^{4}\,{\rm GeV}$ and $10^{9}\,{\rm GeV}$.
+In both Figs.
+3 and 4, we imposed $|\Delta|>100\,{\rm keV}$ and took into account the perturbativity, $|\lambda_{5,{\rm eff}}|<1$.
 Then, we need the singlet scalar mass $m_{S}$ to be of order $10^{14}-10^{16}\,{\rm GeV}$ for $m_{\nu}=0.01-0.1\,{\rm eV}$, which might hint at a connection of neutrino mass generation to the unification.
-If we choose smaller Yukawa couplings $y_{N}$, the diagonal boundary in the left plot of Fig.
-3 is shifted to the left, allowing for smaller $M_{N}$ for a fixed $m_{0}$, and the strip in the right plot of Fig.
+If we choose smaller Yukawa couplings $y_{N}$, the diagonal boundary in Fig.
+3 is shifted to the left, allowing for smaller $M_{N}$ for a fixed $m_{0}$, and the strip in Fig.
 4 is shifted to the bottom, making $m_{S}$ smaller for a fixed $m_{\nu}$.
-Dark matter direct detection: When the lighter neutral scalar from the inert doublet is dark matter, we can avoid the strong bound from direct detection if the mass splitting between $A_{0}$ and $H_{0}$ is large enough, i.e.
+Figure 3: Inert doublet scalar mass $m_{0}$ vs RH neutrino mass $M_{N}$ for varying $\lambda_{5,{\rm eff}}$.
+Figure 4: Neutrino mass $m_{\nu}$ vs singlet scalar mass $m_{S}$ for varying $M_{N}$
+V Dark matter direct detection
+When the lighter neutral scalar from the inert doublet is dark matter, we can avoid the strong bound from direct detection if the mass splitting between $A_{0}$ and $H_{0}$ is large enough, i.e.
 $|\Delta|\gtrsim 100\,{\rm keV}$.
-Then, the cross section for spin-independent scattering between dark matter and nucleus is dominated by the tree-level Higgs exchange Barbieri:2006dq ; LopezHonorez:2006gr , as follows,
-$\sigma^{\rm SI}_{A_{0}-N}=\frac{(\lambda_{L}-\lambda_{5,\,{\rm eff}})^{2}\mu^{2}_{N}}{4\pi m^{2}_{A_{0}}m^{4}_{h}A^{2}}(Zf_{p}+(A-Z)f_{n})^{2}$ (10)
-where $Z,A-Z$ are the numbers of protons and neutrons in the detector nucleus, $\mu_{N}$ is the reduced mass for the DM-nucleus system, and $f_{p,n}$ are the nucleon form factors Hoferichter:2015dsa ; Junnarkar:2013ac ; Kim:2023pwf .
-The DM relic density is determined by the DM annihilations into a pair of SM fermions, $WW,ZZ$ and $h_{\rm SM}h_{\rm SM}$ with $h_{\rm SM}$ being the SM Higgs boson Barbieri:2006dq ; LopezHonorez:2006gr ; Gong:2012ri .
-In particular, there is a room to adjust $\lambda_{L}$ to explain the DM relic density for relatively heavy DM due to $A_{0}A_{0}\to h_{\rm SM}h_{\rm SM}$ while the direct detection bound is satisfied Belyaev:2016lok .
-DM masses above about $100\,{\rm GeV}$ are consistent with the bounds from direct detection, as far as $|\lambda_{L}-\lambda_{5,\,{\rm eff}}|\lesssim 10^{-2}$, apart from the resonance.
-But, if $|\lambda_{L}-\lambda_{5,\,{\rm eff}}|\lesssim 10^{-3}$, as required near resonances, the one-loop corrections to the DM-nucleus cross section are relevant Abe:2015rja .
-Thermal leptogenesis: Baryogenesis can be realized in our model via thermal leptogenesis Fukugita:1986hr , due to the decays of RH neutrinos with hierarchical masses.
+Then, the cross section for spin-independent scattering between dark matter and nucleus is dominated by the tree-level exchanges of Higgs-like scalars Barbieri:2006dq ; LopezHonorez:2006gr , as follows,
+$\sigma^{\rm SI}_{A_{0}}=\frac{\lambda_{L5}^{2}\mu^{2}_{N}(Zf_{p}+(A-Z)f_{n})^{2}}{4\pi m^{2}_{A_{0}}A^{2}}\bigg{(}\frac{c^{2}_{\alpha}}{m^{2}_{h_{1}}}-\frac{s^{2}_{\alpha}}{m^{2}_{h_{2}}}\bigg{)}^{2}$ (10)
+where $\lambda_{L5}\equiv\lambda_{L}-\lambda_{5,\,{\rm eff}}$, $m_{h_{1,2}}$ are the masses for SM Higgs-like and dark Higgs-like scalars, $Z,A-Z$ are the numbers of protons and neutrons in the detector nucleus, $\mu_{N}$ is the reduced mass for the DM-nucleus system, and $f_{p,n}$ are the nucleon form factors Hoferichter:2015dsa ; Junnarkar:2013ac ; Kim:2023pwf .
+The DM relic density is determined by the DM annihilations into a pair of SM fermions, $WW,ZZ$ and $h_{i}h_{j}$ with $i,j=1,2$ Barbieri:2006dq ; LopezHonorez:2006gr ; Gong:2012ri .
+If $H_{0},H^{\pm}$ have comparable masses to $m_{A_{0}}$, we also need to include the co-annihilation channels associated with them as eq.
+(57) in the appendix.
+In particular, there is room to adjust $\lambda_{L}$ and dark Higgs-portal couplings, denoted by ${\cal L}_{\rm int}\supset-\frac{1}{2}\sum_{i=1,2}\lambda_{H_{i}\varphi}|H_{i}|^{2}\varphi^{2}$, to explain the DM relic density for relatively heavy DM due to $A_{0}A_{0}\to h_{i}h_{j}$ while the direct detection bound is satisfied Belyaev:2016lok .
+DM masses above about $100\,{\rm GeV}$ are consistent with the bounds from direct detection, as far as $|\lambda_{L5}(c^{2}_{\alpha}-s^{2}_{\alpha}m^{2}_{h_{1}}/m^{2}_{h_{2}})|\lesssim 10^{-2}$, apart from the resonance, and the bounds from the SM Higgs decays can be easily satisfied for $m_{A_{0}},m_{h_{2}}\gtrsim m_{h_{1}}/2=62.5\,{\rm GeV}$.
+But, for a smaller value of $|\lambda_{L5}|$, as required near resonances, the one-loop corrections to the DM-nucleus cross section are relevant Abe:2015rja .
+In Fig.
+5, we depict the DM relic density as a function of DM mass for $\lambda_{H_{2}\varphi}=0,0.01,0.03$, in red solid, blue dashed and dotted lines, respectively.
+We chose a zero Higgs mixing, $\alpha=0$, the singlet parameters are $\lambda_{H_{1}\varphi}=0$, $v_{\varphi}=10^{4}\,{\rm GeV}$ and $m_{h_{2}}=m_{\rho}=80\,{\rm GeV}$, and the parameters for the inert doublet are $\lambda_{L5}=0.01$, $\Delta=10\,{\rm MeV}$ and $m_{H^{\pm}}=m_{0}+1\,{\rm GeV}$.
+Thus, as $\lambda_{H_{2}\varphi}$ increases, the correct relic density can be explained for a heavier DM mass far beyond $100\,{\rm GeV}$.
+The results are insensitive to $\Delta$ as far as $\Delta\lesssim T_{f}$ where $T_{f}$ is the freeze-out temperature given by $T_{f}\simeq m_{0}/20$.
+Moreover, we remark that a nonzero $\lambda_{H_{1}\varphi}$ also contributes to the dark matter annihilation into $\varphi\varphi$ with the SM Higgs exchange, but its contribution is suppressed for a small $|\lambda_{L5}|$.
+Figure 5: DM relic density as a function of DM mass $m_{0}$, depending on the inert doublet portal coupling, $\lambda_{H_{2}\varphi}$.
+VI Thermal leptogenesis
+Baryogenesis can be realized in our model via thermal leptogenesis Fukugita:1986hr , due to the decays of RH neutrinos with hierarchical masses.
+The RH neutrinos can be once in thermal equilibrium with the SM plasma in the early universe due to their Yukawa couplings to the inert doublet or the spurion scalar, which mixes with the SM Higgs.
 As the asymmetries generated from the decays of the heavier RH neutrinos, $N_{2,3}$, are negligible due to strong washout effects by scattering, we focus on the $N_{1}$ contribution to the lepton asymmetry.
@@
 Here, $\Gamma_{1}$ is the decay rate for $N_{1}$, given by $\Gamma_{1}\simeq\frac{M_{N,1}}{8\pi}(y^{\dagger}_{N}y_{N})_{11}$ for $m_{0}\ll M_{N,1}$, and $H(z_{1})$ is the Hubble parameter.
-In the Casas-Ibarra parametrization for the Yukawa matrix $y_{N}$ Casas:2001sr , we can express $(y^{\dagger}_{N}y_{N})_{ij}=\sqrt{\Lambda_{i}\Lambda_{j}}{\widetilde{m}}_{ij}$ in terms of the effective neutrino mass matrix ${\widetilde{m}}\equiv RD_{{\cal M}_{\nu}}R^{\dagger}$ with $R$ being arbitrary complex rotation matrix, $D_{{\cal M}_{\nu}}={\rm diag}(m_{\nu_{1}},m_{\nu_{2}},m_{\nu_{3}})$ and $\Lambda_{i}=\frac{4\pi^{2}\xi_{i}M_{N,i}}{\lambda_{5,{\rm eff}}v^{2}_{H}}$ Hugle:2018qbw .
+In the Casas-Ibarra parametrization for the Yukawa matrix $y_{N}$ Casas:2001sr , we can express $(y^{\dagger}_{N}y_{N})_{ij}=\sqrt{\Lambda_{i}\Lambda_{j}}{\widetilde{m}}_{ij}$ in terms of the effective neutrino mass matrix ${\widetilde{m}}\equiv RD_{{\cal M}_{\nu}}R^{\dagger}$ with $R$ being arbitrary complex rotation matrix, $D_{{\cal M}_{\nu}}={\rm diag}(m_{\nu_{1}},m_{\nu_{2}},m_{\nu_{3}})$, and $\Lambda_{i}=\frac{4\pi^{2}\xi_{i}M_{N,i}}{\lambda_{5,{\rm eff}}v^{2}_{H}}$ Hugle:2018qbw .
 Here, $\xi_{i}=\big{(}\frac{1}{8}\frac{M^{2}_{N,i}}{m^{2}_{H_{0}}-m^{2}_{A_{0}}}[L(m^{2}_{H_{0}})-L(m^{2}_{A_{0}})]\big{)}^{-1}$ with $L(m^{2})=\frac{m^{2}}{m^{2}-M^{2}_{N,i}}\ln(m^{2}/M^{2}_{N,i})$ are order one parameters in most of the parameter space.
 Thus, we get the decay parameter as $K_{1}\simeq\frac{2\pi^{2}\xi_{1}}{\lambda_{5\,{\rm eff}}}\sqrt{\frac{45}{64\pi^{5}g_{*}}}\frac{M_{P}}{v^{2}_{H}}{\widetilde{m}}_{11}$ for $m_{0}\ll M_{N,1}$.
+Figure 6: Parameter space for RH neutrino mass $M_{N}$ vs $\lambda_{5,{\rm eff}}$, showing $\eta_{B}>6.1\times 10^{-10}$ in purple region.
+The contour lines show the parameter space for the heaviest neutrino mass equal to $m_{\nu}=0.05\,{\rm eV}$.
 For the case with three RH neutrinos, the $CP$ asymmetry parameter is bounded Davidson:2002qv ; Hugle:2018qbw by
@@
 $\eta_{B}\simeq 3.8\times 10^{-10}\Big{(}\frac{M_{N,1}}{10^{4}\,{\rm GeV}}\Big{)}\bigg{(}\frac{10^{-4}}{\lambda_{5,{\rm eff}}}\bigg{)}\bigg{(}\frac{m_{h}}{0.1\,{\rm eV}}\bigg{)}.$ (15)
-Figure 4: Parameter space for RH neutrino mass $M_{N}$ vs $\lambda_{5,{\rm eff}}$, showing $\eta_{B}>6.1\times 10^{-10}$ in purple region.
 […]
@@
 It turns out that $\Delta L=2$ processes are inefficient if $|\lambda_{5,{\rm eff}}|\gtrsim 10^{-5}$ for $M_{N,1}\lesssim 10^{9}\,{\rm GeV}$ Clarke:2015hta .
-Conclusions: We presented a unified solution with $Z_{4}$ discrete Peccei-Quinn symmetry for small neutrino masses and stability of dark matter in the radiative seesaw scenarios.
-Once the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of an extra singlet scalar field $\varphi$ much below the mass scale of the singlet scalar $S$, a small mixing between the dark sector neutral scalars is generated.
-Then, we showed that small neutrino masses are obtained from one-loop diagrams with RH neutrinos and inert doublet scalars with small mass splittings and identified the parameter space where observed neutrino masses are explained and the lighter scalar of the inert doublet becomes a dark matter candidate.
-From the measurement of the absolute neutrino masses at the level of $0.01\,{\rm eV}$, direct detection searches for the inert doublet dark matter and the successful leptogenesis, we showed that a new mass scale of the singlet scalar $S$ is expected to be up to $\sim 10^{16}\,{\rm GeV}$.
+VII Conclusions
+We presented a unified solution with $Z_{4}$ discrete symmetry for small neutrino masses and stability of dark matter in the radiative seesaw scenarios.
+Once the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion singlet scalar $\varphi$ much below the mass of the singlet scalar $S$, a small effective lepton number violating interaction between two Higgs doublets is generated, being responsible for a small mass splitting of the neutral inert scalars at tree level and small neutrino masses at one-loop level.
+We showed that the spurion scalar opens up a new parameter space for the correct relic density for dark matter and a successful leptogenesis is achieved.
 Acknowledgments
 The work is supported in part by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567 and RS-2024-00341419 (JKK)).
+This research was supported by the Chung-Ang University research grant in 2019.

### v4

@@
 The Standard Model is extended with an inert doublet scalar, a dark singlet scalar, a spurion scalar and right-handed neutrinos, which all transform nontrivially under $Z_{4}$.
-After the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion, much below the mass scale of the dark singlet scalar, a small lepton number violating coupling for the inert doublet is generated at tree level, so small neutrino masses are obtained at one-loops.
+After the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion, much below the mass scale of the dark singlet scalar, a small lepton number violating coupling for the inert doublet is generated at tree level, so small neutrino masses are obtained at one-loops for relatively light new fields.
 We discuss the important roles of the $Z_{4}$ symmetry for neutrino masses, dark matter physics and thermal leptogenesis.
@@
 The lepton number is respected in the renormalizable interactions due to $Z_{4}$.
-Then, after the $Z_{4}$ symmetry is broken to $Z_{2}$ due to the VEV of the spurion field, the lepton number is broken, and the lightest neutral dark scalar in the model becomes a dark matter candidate.
+Then, only after the $Z_{4}$ symmetry is broken to $Z_{2}$ due to the VEV of the spurion field, the lepton number is broken, so neutrino masses are generated at one-loops.
+Moreover, the lightest neutral dark scalar in the model becomes a dark matter candidate.
 The spurion scalar can be regarded as a pseudo-Goldstone boson coming from the breaking of a global $U(1)$ symmetry into $Z_{4}$ at a higher scale, and $Z_{4}$ can be further down to $Z_{2}$ at a low scale.
-Taking the decoupling limit for making the singlet scalar field $S$ heavy, we show how a small mass splitting between the neutral components of the inert doublet arises, in turn giving rise to small neutrino masses at one-loops.
+Taking the decoupling limit where the singlet scalar field $S$ becomes heavier than the $Z_{4}$ breaking scale, we show how a small mass splitting between the neutral components of the inert doublet arises, in turn giving rise to small neutrino masses at one-loops.
+In the present case, we argue that large masses of extra fields or small Yukawa couplings are not necessary, unlike the case with the $Z_{2}$ symmetry only.
 We discuss the roles of the new singlet scalars and constraints for neutrino masses, dark matter and thermal leptogenesis.
@@
 (53) in the appendix.
-Then, when the RH neutrinos are heavier than the inert doublet Higgs, namely, $M_{N,k}\gg m_{0}$, the above expression with eq.
+Then, when the RH neutrinos are heavier than the inert doublet Higgs, namely, $M_{N,k}\gtrsim m_{0}$, the above expression with eq.
 (4) gets simplified to
@@
 So, for a small $\lambda_{5,{\rm eff}}$, it is sufficient to take $m_{S}\gtrsim M_{N,k}\sim 10^{9}\,{\rm GeV}$, unless we tune the Yukawa couplings $y_{N,ij}$ to small values.
-In the case with $m_{0}\gg M_{N,k}$, the lightest RH neutrino is a dark matter candidate Ibarra:2016dlb whereas the scalars of the inert doublet decay into a pair of the SM lepton and the RH neutrino through the Yukawa interactions.
+In the case with $m_{0}\gtrsim M_{N,k}$, the lightest RH neutrino is a dark matter candidate Ibarra:2016dlb whereas the scalars of the inert doublet decay into a pair of the SM lepton and the RH neutrino through the Yukawa interactions.
 In this limit, eq.
@@
 Once the $Z_{4}$ symmetry is broken to $Z_{2}$ by the VEV of the spurion singlet scalar $\varphi$ much below the mass of the singlet scalar $S$, a small effective lepton number violating interaction between two Higgs doublets is generated, being responsible for a small mass splitting of the neutral inert scalars at tree level and small neutrino masses at one-loop level.
-We showed that the spurion scalar opens up a new parameter space for the correct relic density for dark matter and a successful leptogenesis is achieved.
+Although the $Z_{2}$ symmetry is sufficient for stability of dark matter, its extension to the $Z_{4}$ symmetry in our model allows for explaining the origin of small neutrino masses, even if inert doublet scalar and RH neutrinos are relatively light and the Yukawa couplings for neutrinos are sizable.
+This option gives rise to phenomenologically a novel window for new particle searches.
+Moreover, we showed that the spurion scalar responsible for the $Z_{4}$ symmetry breaking opens up a new parameter space for obtaining the correct relic density for inert doublet-like dark matter and a successful leptogenesis is achieved.
 Acknowledgments